### v1

# DS-VTON: High-Quality Virtual Try-on via Disentangled Dual-Scale Generation

Abstract
Despite recent progress, most existing virtual try-on methods still struggle to simultaneously address two core challenges: accurately aligning the garment image with the target human body, and preserving fine-grained garment textures and patterns. In this paper, we propose DS-VTON, a dual-scale virtual try-on framework that explicitly disentangles these objectives for more effective modeling. DS-VTON consists of two stages: the first stage generates a low-resolution try-on result to capture the semantic correspondence between garment and body, where reduced detail facilitates robust structural alignment. The second stage introduces a residual-guided diffusion process that reconstructs high-resolution outputs by refining the residual between the two scales, focusing on texture fidelity. In addition, our method adopts a fully mask-free generation paradigm, eliminating reliance on human parsing maps or segmentation masks. By leveraging the semantic priors embedded in pretrained diffusion models, this design more effectively preserves the person’s appearance and geometric consistency. Extensive experiments demonstrate that DS-VTON achieves state-of-the-art performance in both structural alignment and texture preservation across multiple standard virtual try-on benchmarks.
Figure 1: Comparison between existing methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) and our proposed DS-VTON. (a) Existing approaches typically adopt a single-scale pipeline with masked inputs, limiting their ability to capture full-body semantics and garment structure. (b) In contrast, DS-VTON employs a fully mask-free, dual-scale framework: the low-resolution stage focuses on garment-body alignment, while the high-resolution stage refines texture details via residual-guided denoising.
1 Introduction
Given a garment image and a person image, the goal of virtual try-on is to synthesize a photorealistic image of the person wearing the specified garment (Han et al., 2018). As a key enabling technology for online fashion and e-commerce, virtual try-on has attracted increasing attention in recent years (Choi et al., 2021; Ge et al., 2021b; Gou et al., 2023; Lee et al., 2022; Morelli et al., 2022, 2023; Choi et al., 2024; Chong et al., 2025; Zhou et al., 2025). This task involves two fundamental challenges: (1) accurately fitting the garment onto the human body, and (2) preserving fine-grained garment textures. Existing methods fall into two main categories: Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) and Diffusion Models (Ho et al., 2020; Rombach et al., 2022). Early GAN-based approaches (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023) typically follow a two-stage pipeline: a warping module first aligns the garment with the target pose, followed by a generation module to synthesize the final image. While warping helps preserve garment appearance, the subsequent generation stage often leads to detail loss due to imperfect feature fusion.
Recent diffusion-based methods (Kim et al., 2024; Zhu et al., 2023; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025) address both structural alignment and texture preservation within a unified denoising process. Early diffusion steps tend to capture global structure, while later steps refine texture details. This progressive generation order aligns naturally with the requirements of virtual try-on. However, relying on a single-stage diffusion process remains inherently limited. In practice, many existing approaches still struggle to simultaneously ensure accurate garment-body alignment and high-fidelity texture reconstruction. The unified framework may fail to fully disentangle structure from detail, leading to compromised visual quality in the final output.
To address these challenges, we propose a dual-scale framework that explicitly decouples structural alignment and texture refinement. Motivated by the natural progression from structure to detail in diffusion processes, our pipeline first performs coarse alignment and subsequently enhances garment details in a second refinement stage. Specifically, our framework consists of two sequential stages:
  • Low-resolution stage: the model generates a coarse try-on result by suppressing high-frequency content and focusing on accurate structural alignment.
  • High-resolution stage: we extend a residual-guided denoising strategy that predicts the residual between the high-resolution image and its low-resolution counterpart. This enables the model to focus specifically on texture restoration, building upon the structurally aligned base from the first stage.
Beside, traditional virtual try-on methods rely on human parsing masks for spatial guidance, our approach adopts a mask-free strategy that eliminates this dependency by leveraging the strong semantic priors embedded in pretrained diffusion models. This design avoids errors caused by inaccurate masks and preserves the full-body appearance and geometry, allowing the model to adaptively infer garment regions. As a result, our method improves the accuracy of body structure modeling and better retains the individuality of the person’s appearance, leading to more realistic and personalized try-on results. Extensive experiments on VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) validate the effectiveness of our method, demonstrating state-of-the-art performance both qualitatively and quantitatively.
Figure 2: Overview of our proposed method. The overall pipeline consists of a two-scale generation process: the low-resolution stage produces a coarse try-on result, which is then refined by the high-resolution stage. Both stages share the same network architecture. See Section 3 for further details.
2 Related works
GAN-based virtual try-on.
Earlier methods (Choi et al., 2021; Ge et al., 2021a; Lee et al., 2022; Xie et al., 2023), which are based on Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), typically decompose the virtual try-on task into two stages: (1) warping the garment to align with the human body shape, and (2) integrating the warped garment with the human image to generate the final result. For instance, ACGPN (Yang et al., 2020) employs a warping module based on Thin-Plate Spline (TPS) (Duchon, 1977) to deform the garment. PFAFN (Ge et al., 2021b) proposes a parser-free method that guides the garment warping process using learned appearance flows. VITON-HD (Choi et al., 2021) introduces a specialized normalization layer and generator design to better handle garment-body misalignment during synthesis. However, a key limitation of GAN-based approaches is their constrained capacity in capturing both precise spatial alignment and fine-grained garment details. As a result, these methods often rely heavily on the warping stage to encode garment appearance early in the pipeline. Yet, the subsequent integration stage frequently introduces artifacts or detail loss, degrading the overall realism of the try-on result.
Diffusion-based virtual try-on.
With the rapid advancement of diffusion models (Ho et al., 2020; Rombach et al., 2022), a series of powerful virtual try-on methods have been proposed (Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024; Zhou et al., 2025; Sun et al., 2024). Early approaches like DCI-VTON (Gou et al., 2023) follow a traditional two-stage pipeline: first warping the garment to match the body, then applying a diffusion model to blend it with the person image. More recent methods (Morelli et al., 2023; Kim et al., 2024; Zhu et al., 2023; Choi et al., 2024) adopt a single diffusion process that directly synthesizes the try-on result, shifting the focus to more effective conditioning strategies. For example, LaDI-VTON (Morelli et al., 2023) employs textual inversion to encode garment identity, while IDM-VTON (Choi et al., 2024) introduces a GarmentNet module for structural and appearance guidance. Leffa (Zhou et al., 2025) further proposes a Leffa loss to guide attention weights during the final 500 denoising steps, enhancing fine-grained texture recovery. Complementary to these methods, FitDiT (Jiang et al., 2024) explores an alternative direction. Built on SD3 (Esser et al., 2024), which adopts a DiT-based architecture (Peebles and Xie, 2023), it uses an aggressive rectangular mask to impose stronger spatial constraints, effectively mitigating the alignment issues faced by earlier mask-based pipelines. Despite notable progress, current diffusion-based methods still suffer from visible shortcomings. These include garment fragmentation artifacts due to imprecise human segmentation, as well as inaccurate rendering of fine patterns such as flowers or text. Addressing these open challenges remains a key direction.
3 Methodologies
Notations.
Given a person image $\mathbf{x}_{p}\in\mathbb{R}^{H\times W\times 3}$ and a garment image $\mathbf{x}_{g}\in\mathbb{R}^{H\times W\times 3}$, the virtual try-on task aims to generate a realistic output $\mathbf{x}_{r}\in\mathbb{R}^{H\times W\times 3}$, where the person appears to be wearing the given garment. Our pipeline requires no auxiliary conditions such as segmentation masks or human parsing maps. The only additional variables appear in the low-resolution stage, where the inputs $\mathbf{x}_{p}$ and $\mathbf{x}_{g}$ are downsampled to $\tilde{\mathbf{x}}_{p},\tilde{\mathbf{x}}_{g}\in\mathbb{R}^{h\times w\times 3}$, with $h=H/\sigma$, $w=W/\sigma$, and $\sigma$ denoting the downsampling ratio. These are used to generate a low-resolution try-on result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$. For notational simplicity, we assume that $\tilde{\mathbf{x}}_{r}$ has been upsampled to the original resolution when used in the high-resolution stage, and we do not introduce a separate symbol for it. Further details on the low-resolution process and the choice of $\sigma$ are provided in Subsection 3.3.
Mask-free strategy.
Previous virtual try-on methods (Kim et al., 2024; Xu et al., 2025; Choi et al., 2024; Zhou et al., 2025; Jiang et al., 2024) commonly rely on external human parsers to obtain body segmentation masks. These masks are then used to explicitly isolate the garment region in the person image, serving as spatial guidance during the try-on process. While this strategy is reasonable in conventional pipelines, it becomes less compelling in the context of diffusion-based models like Stable Diffusion (Rombach et al., 2022), which are pretrained on large-scale data and inherently encode strong semantic priors related to human structure and clothing layout. Modern pretrained diffusion models already demonstrate robust understanding of body geometry and clothing regions without requiring explicit mask supervision. Motivated by this observation, we adopt a fully mask-free design. Unlike prior work (Ge et al., 2021b), we do not distill pseudo-masks from human parsers. Instead, during both training and inference, the model directly takes a garment image and a person image as input, without relying on any parsing or segmentation guidance. This design offers two major advantages. First, segmentation masks may contain errors, either omitting parts of the garment or incorrectly covering body regions, leading to unrealistic or inconsistent outputs. Second, to avoid garment omission, large masks are sometimes used, which can obscure body regions and impair the model’s ability to reconstruct accurate body shape. In contrast, our mask-free strategy allows the model to adaptively infer the true garment region while preserving full-body geometry. This leads to more faithful body reconstruction and benefits downstream stages, ultimately improving the visual quality of the final try-on result. Please refer to Subsection 4.2 for qualitative comparisons.
3.1 Overview
Our method is built upon Stable Diffusion (Rombach et al., 2022). The backbone architecture follows the dual U-Net framework (Zhang, 2023; Hu, 2024; Xu et al., 2024), which has also demonstrated strong performance in virtual try-on tasks (Choi et al., 2024; Xu et al., 2025). Further architectural details are provided in Subsection 3.2. As shown in Figure 2, both the low-resolution and high-resolution stages share this network architecture.
In the low-resolution stage, we first use a VAE (Kingma, 2013) to encode the downsampled garment image $\tilde{\mathbf{x}}_{g}$, which is then passed to the reference U-Net. The downsampled person image $\tilde{\mathbf{x}}_{p}$ is also VAE-encoded and concatenated along the feature channel with a Gaussian noise tensor of the same shape. This combined latent is fed into the denoising U-Net to generate a low-resolution result $\tilde{\mathbf{x}}_{r}\in\mathbb{R}^{h\times w\times 3}$.
In the high-resolution stage, $\mathbf{x}_{g}$ and $\mathbf{x}_{p}$ are used in the same manner as in the low-resolution stage: $\mathbf{x}_{g}$ is passed to the reference U-Net, while $\mathbf{x}_{p}$ is encoded and concatenated with the latent, which is then input to the denoising U-Net. The main difference is the initialization of the latent. Specifically, instead of initializing $\mathbf{x}_{T}$ as standard Gaussian noise, we define the noisy latent input as:
$\mathbf{x}_{T}=\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r},$ (1)
where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\alpha,\beta$ are balancing coefficients. The model is trained to gradually denoise this latent, transitioning $\mathbf{x}_{T}$ to $\mathbf{x}_{r}$ by converting the noise component $\alpha\cdot\bm{\epsilon}$ into the residual term $\mathbf{x}_{r}-\beta\cdot\tilde{\mathbf{x}}_{r}$, such that the final result satisfies $\mathbf{x}_{0}=\mathbf{x}_{r}$. We detail the full formulation, as well as the design choices for $\alpha$ and $\beta$, in Subsection 3.4.
3.2 Network architecture
We adopt a dual U-Net architecture, where the reference U-Net encodes garment features and integrates them into the main denoising U-Net through self-attention layers. This structure has proven effective for maintaining garment fidelity and enhancing visual quality (Choi et al., 2024), and is retained in our architecture. Following the design philosophy of (Chong et al., 2025), we remove all cross-attention layers in the U-Net, relying solely on self-attention mechanisms. This simplification improves performance and efficiency, as confirmed by our ablation studies (refer to Appendix B.1). All garment-related information is delivered through the reference U-Net, eliminating the need for additional conditioning pathways. To further improve computational efficiency, we execute the reference U-Net only once per sample, treating it purely as a conditioning module, in line with (Li et al., 2024).
We initialize our U-Net weights using those of Stable Diffusion 1.5 (Rombach et al., 2022). While SDXL (Podell et al., 2025) offers more powerful generative capabilities, the goal of our work is to propose a lightweight yet effective framework. Hence, we retain SD1.5 as our backbone. We also conducted preliminary experiments based on transformer architectures such as SD3 and SD3.5, with additional discussion provided in Appendix C. In conclusion, SD1.5 offers an optimal trade-off between model simplicity and try-on performance, and forms the foundation of our pipeline.
3.3 Low-resolution stage
Low-resolution images inherently emphasize structural information due to the suppression of fine-grained textures. Thus, we first introduce a low-resolution stage to guide the overall generation process. This stage serves two main purposes. First, it enables the model to capture the human body shape and garment category, ensuring accurate structural alignment between the person and the clothing. Second, it provides coarse but semantically reliable garment structure, such as the placement of stripes or patterns, offering guidance for fine-detail reconstruction in the high-resolution stage.
In this stage, we first downsample the garment image $\mathbf{x}_{g}$ and the person image $\mathbf{x}_{p}$ by the downsampling ratio $\sigma$, and then perform denoising from pure Gaussian noise $\epsilon$ to generate the low-resolution result $\tilde{\mathbf{x}}_{r}$. Both training and inference follow the standard diffusion process used in Stable Diffusion (Rombach et al., 2022; Ho et al., 2020; Song et al., 2020).
The only hyperparameter in this stage is the downsampling ratio $\sigma$. We consider three values: $\sigma=1,2,$ and $4$. When $\sigma=1$, the low-resolution stage operates at the same resolution as the high-resolution stage; $\sigma=2$ corresponds to downsampling from $768\times 1024$ to $384\times 512$, and so on. Based on our observations, downsampling ratios of $\sigma=2$ and $\sigma=4$ both enhance human body structure understanding, while $\sigma=2$ offers the most reliable structural guidance for accurate garment reconstruction. Quantitative and qualitative comparisons of different $\sigma$ values are presented in Subsection 4.3, and $\sigma=2$ is used in all experiments throughout this work.
3.4 High-resolution stage
While the low-resolution stage focuses on body structure by ignoring fine details, the high-resolution stage is designed to refine garment textures. Instead of directly reconstructing the full high-resolution image, this stage aims to reconstruct the residual between the high- and low-resolution outputs. This strategy effectively disentangles structural alignment from detailed appearance modeling.
3.4.1 Reformulating the denoising process with residual learning
DDPM.
Denoising Diffusion Probabilistic Mode (DDPM) (Ho et al., 2020) aims to approximate the true data distribution by leveraging the diffusion probabilistic model framework (Sohl-Dickstein et al., 2015), which defines a Markov chain of length $T$ that gradually transforms pure Gaussian noise into a sample from the data distribution. Compared with earlier diffusion models, DDPM incorporates ideas from score matching (Song and Ermon, 2019), simplifying the objective by training the model to predict only the noise component $\bm{\epsilon}$, which approximates the score function (i.e., the gradient of the log-density). The forward and reverse diffusion processes are defined as:
$\mathbf{x}_{t}$ $=\sqrt{\alpha_{t}}\,\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon}$
$=\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon},$ (2)
$\mathbf{x}_{t-1}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}.$ (3)
Equations (2) and (3) define the forward and reverse diffusion processes, respectively. In these equations, $\bm{\epsilon},\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denotes the noise predicted by the model. The parameters $\alpha_{t}$ and $\bar{\alpha}_{t}$ are predefined noise scheduling terms. In this framework, $\mathbf{x}_{0}\sim p_{data}(\mathbf{x}_{0})$ denotes a sample from the true data distribution, while $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. All formulations above hold for $t=1,2,\dots,T$.
Residual-guided denoising reformulation.
Inspired by  (Yue et al., 2023), we reformulate the forward and reverse processes to explicitly incorporate the low-resolution generation result $\tilde{\mathbf{x}}_{r}$ as a structural prior for the high-resolution generation. Instead of constructing a Markov chain solely between Gaussian noise and $p_{data}(\mathbf{x}_{0})$, we aim to build a transition path from $\tilde{\mathbf{x}}_{r}$ to the high-resolution result $\mathbf{x}_{r}$. A naive formulation such as $\mathbf{x}_{T}=\bm{\epsilon}+\tilde{\mathbf{x}}_{r}$, $\mathbf{x}_{0}=\mathbf{x}_{r}$ lacks flexibility in controlling the influence of $\tilde{\mathbf{x}}_{r}$ and $\bm{\epsilon}$. To address this, we adopt a more flexible initialization as defined in Equation (1), with $\mathbf{x}_{0}=\mathbf{x}_{r}$. Under this formulation, the residual-guided forward and reverse diffusion processes become:
$\mathbf{x}_{t}$ $=\sqrt{\alpha_{t}}\,\mathbf{x}_{t-1}+\sqrt{1-\alpha_{t}}\left(\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}\right)$
$=\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\left(\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}\right),$ (4)
$\mathbf{x}_{t-1}$ $=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z}.$ (5)
Equations (4) and (5) define the forward and reverse processes, respectively. The model is trained to predict $\tilde{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},t)\approx\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, such that the noise component in $\mathbf{x}_{t}$ is replaced by a composition of $\bm{\epsilon}$ and the low-resolution result $\tilde{\mathbf{x}}_{r}$. Except for this reformulated initialization and noise target, the rest of the denoising process remains consistent with the original DDPM. All formulations above hold for $t=1,2,\dots,T$.
3.4.2 Controlling noise-structure balance
In the high-resolution stage, generation begins from the latent $\mathbf{x}_{T}=\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$, which combines a stochastic noise term and a structural prior. Importantly, although the noise term has been modified, the actual source of randomness remains from $\bm{\epsilon}$. During the reverse denoising process, this component $\alpha\cdot\bm{\epsilon}$ is gradually transformed into the residual $\mathbf{x}_{r}-\beta\cdot\tilde{\mathbf{x}}_{r}$, ensuring that the final output satisfies $\mathbf{x}_{0}=\mathbf{x}_{r}$. This formulation offers explicit control over the trade-off between randomness and structural guidance. A larger $\beta$ may cause the generation to overly rely on the low-resolution input, which is intended to provide only structural guidance, thereby potentially suppressing the recovery of fine details. While a larger $\alpha$ introduces stronger randomness that may override structural consistency. Through empirical studies, we find that setting $\alpha=\beta=0.5$ offers a good trade-off between fidelity and flexibility. More discussions can be found in Section 4.3.
Figure 3: Qualitative comparison on the VITON-HD dataset. DS-VTON(LR) denotes the low-resolution result, and DS-VTON(HR) represents the final high-resolution result.
4 Experiments
4.1 Experimental setup
Datasets.
In this paper, we conduct experiments on the VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. All ablation studies are carried out on VITON-HD. While VITON-HD contains only upper-body garments, DressCode includes three garment categories: upper-body, lower-body, and dresses. Both datasets consist of paired images, each containing a person image and a corresponding garment image. However, as our method is mask-free, the original paired data alone is insufficient for training. To address this, we use IDM-VTON (Choi et al., 2024) to synthesize additional training data. Further details are provided in Appendix A.1.
Implementation details.
For network initialization, both the reference U-Net and the denoising U-Net are initialized with pretrained weights from Stable Diffusion 1.5 (Rombach et al., 2022). As detailed in Subsections 3.3 and 3.4, we set the downsampling ratio $\sigma=2$ and use $\alpha=\beta=0.5$ to initialize $\mathbf{x}_{T}$ in the high-resolution stage. Consequently, the low-resolution stage operates at a resolution of $384\times 512$, while the high-resolution stage produces outputs at $768\times 1024$. During inference, the two stages are executed sequentially with 20 sampling steps each, using the DDIM sampler (Song et al., 2020). Additional training details are provided in Appendix A.2.
Baselines.
We compare our method with several recent state-of-the-art approaches, including CatVTON (Chong et al., 2025), IDM-VTON (Choi et al., 2024), Leffa (Zhou et al., 2025), OOTDiffusion (Xu et al., 2025), and FitDiT (Jiang et al., 2024), using their official model weights and inference code. For fairness, we standardize the number of inference steps to 30 across all methods. All methods, except FitDiT, are trained separately on each dataset and evaluated accordingly. FitDiT provides only a single set of pretrained weights, jointly trained on VITON-HD, DressCode, and CVDD (Jiang et al., 2024), and is included in our evaluation for completeness.
Evaluation metrics.
Previous virtual try-on methods typically evaluate performance under both paired and unpaired settings. The paired setting involves reconstructing the original person image with the same garment, while the unpaired setting replaces it with a different one (Choi et al., 2021). As most prior approaches rely on masking the garment region, they support both settings. In contrast, our method is mask-free and is therefore evaluated only under the unpaired setting, which better reflects real-world scenarios. Following prior works (Chong et al., 2025; Jiang et al., 2024; Zhou et al., 2025), we adopt Fréchet Inception Distance (FID) (Parmar et al., 2022) and Kernel Inception Distance (KID) (Bińkowski et al., 2018) as quantitative metrics. We also conduct a user study to assess perceptual quality: for VITON-HD, we randomly sample 100 results per method; for DressCode, we sample 33, 33, and 34 results from the dresses, lower-body, and upper-body categories, respectively. Participants are asked to select the result they think performs better in the try-on task. All evaluations are conducted at a resolution of $768\times 1024$.
Table 1: Quantitative comparisons on the VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. FitDiT (Jiang et al., 2024), trained jointly on VITON-HD, DressCode, and CVDD (Jiang et al., 2024), is included for completeness. In contrast, all other methods are trained individually on each dataset.
Dataset VITON-HD DressCode
Method FID ↓ KID ↓ User Study ↑ FID ↓ KID ↓ User Study ↑
OOTDiffusion (Xu et al., 2025) 9.02 0.63 4.1 7.10 2.28 7.2
IDM-VTON (Choi et al., 2024) 9.10 1.06 11.6 5.51 1.42 9.1
CatVTON (Chong et al., 2025) 9.40 1.27 3.4 5.24 1.21 5.2
Leffa (Zhou et al., 2025) 9.38 0.92 4.7 6.17 1.90 7.5
FitDiT (Jiang et al., 2024) 9.33 0.89 19.7 4.47 0.41 34.3
DS-VTON (ours) 8.24 0.31 56.5 4.21 0.34 36.7
Table 2: Ablation on dual-scale design and downsampling ratio $\sigma$.
Version FID $\downarrow$ KID $\downarrow$
$\sigma=1$ 8.97 1.01
$\sigma=1,\alpha=\beta=\tfrac{1}{2}$ 8.77 0.61
$\sigma=4,\alpha=\beta=\tfrac{1}{2}$ 8.41 0.57
$\sigma=2,\alpha=\beta=\tfrac{1}{2}$ 8.24 0.31
Table 3: Ablation on coefficients $\alpha,\beta$ under fixed $\sigma=2$.
Version FID $\downarrow$ KID $\downarrow$
$\sigma=2,\alpha=\beta=\tfrac{1}{2}$ 8.24 0.31
$\sigma=2,\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$ 8.46 0.55
$\sigma=2,\alpha=\tfrac{1}{3},\beta=\tfrac{2}{3}$ 8.26 0.35
$\sigma=2,\alpha=\beta=1$ 8.75 0.94
4.2 Quantitative and qualitative results
Qualitative comparison.
Figure 3 presents a qualitative comparison between DS-VTON and recent baseline methods on the VITON-HD (Choi et al., 2021) dataset. Row numbers referenced below correspond to Figure 3. We evaluate each method in terms of two key aspects: structural alignment and detail preservation. In terms of structural alignment, most mask-based methods fail to accurately capture body pose and garment shape, as illustrated in Row 1. FitDiT (Jiang et al., 2024), which uses a larger rectangular mask, performs relatively better but introduces noticeable artifacts: it fails to reconstruct the hands and exhibits visible artifacts at the junction of the upper and lower garments in Row 1. OOTDiffusion (Xu et al., 2025) alters the original skin tone in Row 2. CatVTON (Chong et al., 2025), IDM-VTON (Choi et al., 2024), and Leffa (Zhou et al., 2025) also show varying degrees of misalignment. In contrast, DS-VTON consistently achieves accurate alignment across a wide range of poses and garment types. Regarding detail preservation, CatVTON and IDM-VTON fail to retain fine-grained textures on complex garments (Row 3), with IDM-VTON generating oversimplified or missing patterns. While Leffa, OOTDiffusion, and FitDiT better preserve textures, they each show limitations: FitDiT achieves texture preservation at the cost of distorting the person’s actual body shape and dressing structure (Rows 2 and 4), OOTDiffusion introduces artifact patterns (Row 5), and Leffa exhibits reduced pattern clarity in complex regions (Rows 3). Furthermore, both FitDiT and Leffa suffer from tonal inconsistencies—FitDiT produces noticeably brighter garments (Row 6), while Leffa tends to generate darker outputs. In contrast, DS-VTON preserves fine textures while maintaining tonal fidelity throughout.
Quantitative comparison.
We conduct experiments on both the VITON-HD (Choi et al., 2021) and DressCode (Morelli et al., 2022) datasets. As shown in Table 1, DS-VTON achieves substantial improvements across both benchmarks. CatVTON (Chong et al., 2025) generates images at $384\times 512$, which we upsample to $768\times 1024$ for comparison. To confirm that this does not bias results, we also evaluate it at native resolution, obtaining FID and KID scores of 9.36 and 1.19, respectively—indicating minimal degradation due to upsampling. Unlike prior methods (Chong et al., 2025; Zhou et al., 2025; Choi et al., 2024) that rely on explicit mask generation and inpainting, DS-VTON is entirely mask-free. This enables accurate, high-quality outputs that are robust and reproducible without dependence on mask quality.
4.3 Ablation study
Figure 4: Visualized results under varying downsampling ratios $\sigma$. For each $\sigma$, the LR column shows the output from the low-resolution stage, and the corresponding HR column shows the high-resolution result refined from that output.
Figure 5: Visualized results under different $\mathbf{x}_{T}$ initialization settings ($\mathbf{x}_{T}=\alpha\cdot\bm{\epsilon}+\beta\cdot\tilde{\mathbf{x}}_{r}$).
Ablation on dual-scale design.
To validate the effectiveness of our dual-scale design, we train four variants on the VITON-HD dataset. As shown in Table 3, Row 1 ($\sigma=1$) corresponds to training the model directly at high resolution ($768\times 1024$) without a refinement stage, aligning with the $\sigma=1$ (LR) column in Figure 5. Applying the mask-free strategy at this resolution leads to poor structural results, highlighting the need for coarse-level guidance. Row 2 ($\sigma=1,\alpha=\beta=\frac{1}{2}$), shown in the $\sigma=1$ (HR) column, adds a refinement stage. While some structural errors are alleviated, relying solely on the second stage to recover both structure and detail still results in failures, due to the lack of reliable low-resolution guidance. This also violates our design principle of decoupling structure modeling (stage one) from detail refinement (stage two). Rows 3 and 4 evaluate the full dual-scale pipeline with $\sigma=4$ and $\sigma=2$, respectively. When garments are structurally simple, both perform reasonably well. However, as shown in Row 1 of Figure 5, $\sigma=4$ introduces visible information loss in complex cases. These qualitative results align with the quantitative trends: among the four variants—single-stage ($\sigma=1$) and dual-stage ($\sigma=1$, $\sigma=2$, $\sigma=4$)—the single-stage model performs worst, $\sigma=4$ shows moderate improvement, and $\sigma=2$ achieves the best overall performance.
Ablation on coefficient-based initialization of $\mathbf{x}_{T}$.
When fixing the downsampling ratio to $\sigma=2$, the coefficients $\alpha$ and $\beta$ control how much the low-resolution output influences the initialization of the high-resolution stage. As shown in Figure 5, setting $\alpha$ too high—such as $\alpha=1,\beta=1$ or $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$—leads to structural distortions in the final result, despite the low-resolution output already providing accurate guidance. In both cases, the red stripe on the garment appears distorted, and in the $\alpha=\tfrac{2}{3},\beta=\tfrac{1}{3}$ configuration, the "GANT" text also becomes visibly warped. Among the remaining configurations, $\alpha=\tfrac{1}{3},\beta=\tfrac{2}{3}$ and $\alpha=\tfrac{1}{2},\beta=\tfrac{1}{2}$ yield generally good results. However, the lower-$\alpha$ version ($\alpha=\tfrac{1}{3}$) still shows minor distortion in the red stripe, indicating insufficient control over fine structure. Balancing these observations, we adopt $\alpha=\beta=\tfrac{1}{2}$ as our default setting. The quantitative results in Table 3 are consistent with this choice.
5 Conclusions
We propose DS-VTON, a dual-scale virtual try-on framework that introduces residual-guided denoising to bridge low- and high-resolution generation. This design enables a principled decoupling of the two core challenges in virtual try-on: structural alignment and detail refinement. Combined with a mask-free strategy, our approach achieves significant improvements over existing methods in both visual quality and robustness. Moreover, the proposed low-to-high resolution paradigm is inherently scalable and generalizable, offering potential for extension to higher resolutions and broader generation tasks beyond virtual try-on, such as personalized image synthesis.
Appendix
Appendix A Experimental details
A.1 Additional training data generation and associated challenges
Our method adopts a mask-free paradigm, in contrast to prior approaches [Choi et al., 2024, Xu et al., 2025, Li et al., 2024], which rely on paired person-garment images by masking out the garment region in the person image and reconstructing it using the standalone garment image. In our case, the person image remains unaltered throughout the training process. To train the low-resolution stage, each training sample requires three images: (1) a garment image, (2) a person image wearing that garment, and (3) another person image of the same identity but wearing a different garment. The third image is constructed by randomly sampling another garment of the same category from the dataset and synthesizing a new person image using IDM-VTON [Choi et al., 2024]. For the high-resolution stage, we additionally require the low-resolution output corresponding to the original person-garment pair. To obtain this, we use our trained low-resolution model to generate a coarse try-on result by inputting the target garment and the synthesized person image (i.e., with a different garment). This output is then used as the low-resolution input for the high-resolution stage.
In summary, each high-resolution training sample consists of: (1) the target garment image, (2) a person image wearing a different garment, (3) the corresponding low-resolution try-on result for the target garment, and (4) the ground-truth high-resolution image of the person wearing the target garment.
While the above construction enables mask-free supervision, it inevitably introduces certain artifacts. Specifically, in the low-resolution stage, the reference U-Net encodes the garment image, while the denoising U-Net operates on the concatenation of the noisy latent and the person image—here, a synthesized image of the same identity wearing a different garment.
Since the synthesized person image is generated by a model such as IDM-VTON, it may exhibit variations beyond the garment itself, including changes in hairstyle, background, or the presence of accessories. As a result, the model may inadvertently learn to alter these unrelated regions during training. Ideally, such issues would not occur if fully disentangled and clean ground-truth data were available. Fortunately, we find this side effect to be limited in practice, as most synthesized person images remain visually consistent with the original identity.
A.2 More implementation details
All experiments are conducted on 8 NVIDIA A6000 GPUs. For both VITON-HD [Choi et al., 2021] and DressCode [Morelli et al., 2022], the low-resolution stage is trained with a batch size of 8, and the high-resolution stage with a batch size of 2. Both stages are optimized using the AdamW optimizer [Loshchilov et al., 2017] with a learning rate of $1\text{e}{-6}$. For VITON-HD, the low-resolution and high-resolution stages are trained for 15,000 and 30,000 steps, respectively (approximately 5 and 24 hours). For DressCode, we jointly train all three garment categories, with the two stages trained for 20,000 and 30,000 steps (approximately 7 and 24 hours). During inference, both stages use 20 DDIM sampling steps. On a single A6000 GPU, the low-resolution stage takes about 1 second per sample, and the high-resolution stage about 4 seconds.
Appendix B More ablation studies
B.1 Ablation on removing cross-attention layers in U-Net
Several previous try-on methods [Choi et al., 2024, Xu et al., 2025, Li et al., 2024] based on dual U-Net architectures incorporate additional conditional encoders, such as CLIP [Radford et al., 2021] or IP-Adapter [Ye et al., 2023], to inject garment information via cross-attention. However, these approaches introduce extra complexity, and it remains unclear whether they actually improve performance. To investigate this, we conduct an ablation study on the VITON-HD [Choi et al., 2021] dataset by comparing two versions of our low-resolution stage ($384\times 512$): (1) our baseline model, which does not include any cross-attention layers, and (2) a variant built on the baseline by adding cross-attention layers to both the reference U-Net and the denoising U-Net. In the latter, garment features are encoded using CLIP and injected via cross-attention. As shown in Figure 6 and Table 4, adding cross-attention does not help preserve garment details; on the contrary, it introduces noticeable distortions. This is also reflected in the performance metrics.
Figure 6: Comparison between our baseline architecture and the variant with cross-attention. In the variant, garment features encoded by CLIP are injected into both the reference U-Net and the denoising U-Net via cross-attention layers.
Table 4: Ablation study on the effect of cross-attention layers in U-Net.
Version FID $\downarrow$ KID $\downarrow$
With cross-attention 9.07 0.94
DS-VTON (low-resolution stage) 8.88 0.72
Appendix C Discussions
How about utilizing the DiT architecture for mask-free try-on?
We also explored using the DiT [Peebles and Xie, 2023] architecture for mask-free try-on. Specifically, we implemented two variants based on SD3 and SD3.5 [Esser et al., 2024], constructing a dual-DiT structure analogous to the dual U-Net design. The reference DiT encodes the garment image, while the person image is concatenated with the latent tensor along the sequence dimension and passed to the denoising DiT. To integrate the two branches, we follow a structure-aligned fusion strategy: since the reference DiT and denoising DiT share the same architecture, we concatenate the latent features from the corresponding transformer block of the reference DiT into the denoising DiT block, specifically on the key and value (K, V) inputs of the attention layer, before computing self-attention. We also remove the text encoder input entirely, so the joint attention layers in DiT degenerate into pure self-attention. However, both versions failed to converge during training. We speculate that directly applying DiT to the mask-free try-on task may be suboptimal, and therefore did not pursue further investigation.
How about adopting a refinement mechanism similar to that of SDXL?
In addition to our proposed refinement approach, we also explored an alternative refinement strategy inspired by SDXL [Podell et al., 2025]. In SDXL, the refinement process involves training a separate model that focuses specifically on denoising the final 200 steps of the diffusion process. During inference, noise corresponding to timestep 200 is added to the output of the first-stage model, and the refinement model is then used to denoise and generate the final result. We experimented with two variations of this approach. In the first variant, the first-stage and refinement-stage outputs share the same resolution. However, as shown in Subsection 4.3, directly applying a mask-free strategy at high resolution inevitably leads to structural distortions, which cannot be effectively corrected by the refinement stage. This led us to the second variant, which we refer to as the noising-denoising strategy. Here, a low-resolution result is first generated and then upsampled to high resolution. Noise is added at this stage (corresponding to a specific timestep), and the refinement model performs denoising from that point onward. While this method can sometimes produce visually plausible results when the target patterns are simple, as shown in Table 5 and Figure 7, its overall performance remains suboptimal. The reason, we believe, lies in the difference between data distributions involved. Although our residual-guided denoising strategy also begins with adding noise to the low-resolution output, it differs fundamentally in how it relates the two stages. In the noising-denoising approach, the refinement model learns to denoise samples drawn solely from the high-resolution distribution, and it lacks an explicit mechanism to relate this with the low-resolution inputs. In contrast, our residual-guided denoising bridges the gap between the low-resolution and high-resolution distributions. This connection is key: it allows the model to better capture the transformation between the two distributions involved. Importantly, while the original diffusion process constructs a mapping from a simple, tractable distribution (such as Gaussian noise) to a complex data distribution, the residual-guided refinement builds a direct bridge between two complex distributions.
Figure 7: Comparison between our baseline and the noising-denoising variant.
Table 5: Comparison of different refinement strategies on the VITON-HD dataset.
Version FID $\downarrow$ KID $\downarrow$
Noising-denoising way 8.98 0.65
DS-VTON 8.24 0.31
Broader impacts.
The ability of DS-VTON to generate realistic virtual try-on results at multiple resolutions makes it well-suited for practical deployment in e-commerce scenarios, where different resolutions are often required across platforms and devices. At the same time, as with other generative technologies, DS-VTON may raise concerns related to intellectual property and personal privacy. We encourage its responsible and ethical use.
Limitation and Future Work.
As discussed in Subsection A.1, one key limitation of our method lies in the data generation process. Due to the reliance on synthesized person images, the model may inadvertently learn to alter regions unrelated to the garment (e.g., hair, accessories, or background). While this issue is not severe in most cases, we acknowledge it as a limitation and consider improving data disentanglement and identity preservation an important direction for future work. Another limitation stems from the use of fixed coefficients $\alpha$ and $\beta$ to initialize the high-resolution refinement stage. Although this static strategy proves effective, it may be overly rigid. In future work, we plan to investigate adaptive or learnable coefficient scheduling mechanisms, which could offer more flexible and content-aware refinement during generation.
Appendix D More experiment results
In this section, we present additional qualitative comparison results on the DressCode [Morelli et al., 2022] dataset and more comparison results on the VITON-HD [Choi et al., 2021] dataset.
Figure 8: More qualitative comparison on the VITON-HD dataset. DS-VTON (LR) denotes the low-resolution output, and DS-VTON (HR) represents the final high-resolution result.
Figure 9: Qualitative comparison on the DressCode dataset (Dresses category). DS-VTON (LR) denotes the low-resolution output, and DS-VTON (HR) represents the final high-resolution result.
Figure 10: Qualitative comparison on the DressCode dataset (Lower category). DS-VTON (LR) denotes the low-resolution output, and DS-VTON (HR) represents the final high-resolution result.
Figure 11: Qualitative comparison on the DressCode dataset (Upper category). DS-VTON (LR) denotes the low-resolution output, and DS-VTON (HR) represents the final high-resolution result.